# Robust variational neural posterior estimation for simulation-based inference

### v1

@@ Appendix A Appendix
A.1 Misspecified Observations Overlaying Training Simulations
@@
A.2 RVNP Training Pipeline
@@
1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalizing flow $p_{\Psi}$
@@
1:Input: Observed dataset $\mathcal{O}$, normalizing flow $p_{\Psi}$, posterior $p_{\phi}$, error model network $\bm{\xi}_{\alpha}$
@@
5:   Sample $\bm{\theta}^{(l)}\sim p_{\phi}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
@@
$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big{[}p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big{]}\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{p_{\phi}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
@@
1:Input: Simulator dataset $D$, posterior $p_{\phi}$, error model $\bm{\xi}_{\alpha}$ (fixed)
@@
$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log p_{\phi}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
@@
A.3 Training Procedure
In every task, RVNP is defined using a variational posterior model based on a rational quadratic spline (RQS) flow with $B=10$ and $15$ knots. The depth of the flow is set to $5$ layers, while the neural network conditioner has a hidden block dimension of $52$. The flow dimension is the dimension of $\theta_{\text{dim}}$, and the conditioning dimension corresponds to that of $x_{\text{sim}}$. The simulator flow has the same architecture as the input and output dimensions swapped.
The importance weighted autoencoder objective was trained with a batch size of $1024$ over $500$ iterations, using the Adam optimizer with learning rate $10^{-3}$, momentum term $\beta_{1}=0.9$, $\epsilon=10^{-8}$, weight decay $10^{-5}$, gradient clipping at $10.0$, and a warmup schedule of $1000$ steps. Early stopping is applied with a patience of $100$ iterations, and $10\%$ of the data is reserved for validation. In the forward modelling of the posterior and the simulator, $K_{\text{obs samples}}=30$ is the number of samples used in the importance weighting. In each task, the simulator was trained using the same optimiser parameters but using the maximum likelihood loss.
Neural Statistic Estimator To train the neural statistic estimator on the InfoMax objective, we adopt a neural statistic and a discriminator model. We use the same optimiser parameters and batch size for the main training routine. All models are implemented in Equinox and trained using JAX. The encoder outputs a deterministic latent representation $z$ without variational sampling, and the discriminator maximises the mutual information between $z$ and $\bm{\theta}$. The spectra encoder uses one-dimensional convolutional feature extraction and global attention modelling using a Conformer block. The hidden dimension for both the embeddings and the discriminator is 100. We describe the algorithm in 5.
@@ A.3 Training Procedure
InfoMax Loss Function
To train the encoders, we maximise the mutual information (MI) between the latent embeddings $z$ and the conditioning variables $\theta$. Algorithm 8 summarises the loss function.
Algorithm 8 InfoMax (Shannon) Loss Computation

### v2

@@ Acknowledgments
MOC is supported by the Gianna Angelopoulos Programme for Science, Technology and Innovation (GAPSTI) through the Science and Technology Facilities Council (STFC) studentship for astronomy. GG acknowledges support from The Leverhulme Trust, through Emeritus Fellowship 2025-007. KSM is supported by the European Union’s Horizon 2020 research and innovation programme under European Research Council Grant Agreement No 101002652 (BayeSN) and Marie Skłodowska-Curie Grant Agreement No 873089 (ASTROSTAT-II).
@@ Appendix A Appendix
In this section, we discuss some results that are not central to the main claims of the paper.
A.1 Sample importance resampling for Spectra task well-specified
Sample-importance-resampling is a useful technique in variational autoencoder frameworks, as it adjusts the variational posterior samples based on the learnt forward model. In Figure 7, we show an example of the effect of sample-importance-resampling in the spectra task. From this plot, it is clear that the samples from the variational posterior (green) are sampled in regions of parameter space with extremely low prior probability. The sample-importance-resampling has the effect of tightening the posterior by removing points which are incompatible with the prior and the simulator model.
Figure 7: (SIR) indicates sample-importance-resampling. We show the effect of sample-importance-resampling on the well-specified spectra task. The effect largely reduces the number of points which have very low prior probability and are incompatible with the simulator likelihood proxy.
A.2 Well-specified case
In this subsection, we present plots of the metrics when we use RVNP to infer the posterior distribution when the model is well-specified. We find that RVNP is not significantly detrimental to posterior recovery in the well-specified case, as it will become under-confident and the sample-importance-resampling distribution becomes very close to the NPE estimate. RVNP is quite under-confident in the spectra and pendulum task. However, this is highly preferable to overconfidence. Moreover, the NPE task is highly over-fit in both of these tasks, and even a small perturbation in output space will give highly wrong posteriors. Such a tight posterior would rarely be used in practice.
Figure 8: Results for the well-specified CS task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference when the simulator is well specified. However, the error introduced by the error model will inflate the posterior. Sample-importance-resampling will sample closest to the NPE estimate. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling.
Results.
Figure 9: Results for the well-specified SIR task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference when the model is well-specified. However, the error introduced by the error model will inflate the posterior. Sample-importance-resampling will sample closest to the NPE estimate, and hybrid sample importance-resampling matches the NRMSE of the NPE. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling and should not be confused with the SIR Task.
Results.
Figure 10: Results for the well-specified Pendulum task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference when the model is well-specified. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling
Results.
Figure 11: Results for the well-specified spectra task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling
A.3 Learnt Error Model for $N_{\rm obs}=1000$ in final two summary statistics
We display the final two dimensions of the output of the simulator in each of the tasks.
Figure 12: The final two dimensions of the output of the simulator in the CS task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace. After clipping the lengths of the vectors above a magnitude of 1 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
Figure 13: The final two dimensions of the output of the simulator in the SIR task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace. After clipping the lengths of the vectors above a magnitude of 5 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
Figure 14: The final two dimensions of the output of the simulator in the SIR task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace. After clipping the lengths of the vectors above a magnitude of 1 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
Figure 15: The final two dimensions of the output of the simulator in the SIR task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace. After clipping the lengths of the vectors above a magnitude of 1 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
A.4 Misspecified Observations Overlaying Training Simulations
@@
A.5 RVNP Training Pipeline
@@
1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalising flow $p_{\Psi}$
@@
1:Input: Observed dataset $\mathcal{O}$, normalising flow $p_{\Psi}$, posterior $q_{\bm{\phi}}$, error model network $\bm{\xi}_{\alpha}$
@@
5:   Sample $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
@@
$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big[p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big]\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
@@
1:Input: Simulator dataset $D$, posterior $q_{\bm{\phi}}$, error model $\bm{\xi}_{\alpha}$ (fixed)
@@
$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
@@
A.6 Training Procedure
In every task, RVNP is defined using a variational posterior model based on a rational quadratic spline (RQS) flow with $B=10$ and $15$ knots using the masked autoregressive flow architecture template of papamakarios2018maskedautoregressiveflowdensity and implemented using the flowjax python package. The depth of the flow is set to $5$ layers, while the neural network conditioner has a hidden block dimension of $52$. The flow dimension is the dimension of $\theta_{\text{dim}}$, and the conditioning dimension corresponds to that of $x_{\text{sim}}$. The simulator flow has the same architecture as the input and output dimensions swapped.
The importance weighted autoencoder objective was trained with a batch size of $1024$ over $500$ iterations, using the Adam optimiser with a learning rate of $10^{-3}$, momentum term $\beta_{1}=0.9$, $\epsilon=10^{-8}$, weight decay $10^{-5}$, gradient clipping at $10.0$, and a warmup schedule of $1000$ steps. Early stopping is applied with a patience of $100$ iterations, and $10\%$ of the data is reserved for validation. In the forward modelling of the posterior and the simulator, $K_{\text{obs samples}}=30$ is the number of samples used in the importance weighting. In each task, the simulator was trained using the same optimiser parameters but using the maximum likelihood loss. When sampling with sampling-importance-resampling, we use $5000$ particles per observation and sample $100$ times per particle.
Neural Statistic Estimator To train the neural statistic estimator on the InfoMax objective, we adopt a neural statistic and a discriminator model. We use the same optimiser parameters and batch size for the main training routine. All models are implemented in Equinox and trained using JAX. The encoder outputs a deterministic latent representation $z$ without variational sampling, and the discriminator maximises the mutual information between $z$ and $\bm{\theta}$. The spectra encoder uses one-dimensional convolutional feature extraction and global attention modelling using a Conformer block. The hidden dimension for both the embeddings and the discriminator is 100. We describe the algorithm in 6.
@@ A.6 Training Procedure
InfoMax Loss Function chen2021neuralapproximatesufficientstatistics show that finding an embedding $\iota_{\mathbf{\omega}}$ that maximizes the mutual information between $\mathbf{\theta}$ and $\iota_{\mathbf{\omega}}(\mathbf{x})$ tends towards sufficient statistics of $\mathbf{\theta}$ provided the dimension of the embedding space is not too small. Furthermore, hjelm2018learning introduce the InfoMax objective for learning deep representations of high-dimensional data based on maximizing the mutual information. After adopting a discriminator network $D_{\mathbf{\gamma}}$, we can generate a lower bound on the mutual information using the Shannon-Jensen divergence
$\mathcal{I}_{\mathrm{JSD}}({\omega,\gamma}):=\mathbb{E}_{P}\big[-\operatorname{sp}\big(-D_{\mathbf{\gamma}}(\mathbf{\theta},\iota_{\mathbf{\omega}}(\mathbf{x}))\big)\big]-\mathbb{E}_{P\times\tilde{P}}\big[\operatorname{sp}\big(D_{\mathbf{\gamma}}(\mathbf{\theta},\iota_{\mathbf{\omega}}(\mathbf{x}))\big)\big],$ (23)
where $\operatorname{sp}(u)=\log(1+e^{u})$ is the softplus function, $P$ is the joint distribution of $(\mathbf{\theta},\iota_{\mathbf{\omega}}(\mathbf{x}))$, and $\tilde{P}$ denotes the product of marginals. Jointly optimizing for the embedding and the discriminator maximises a lower bound on the mutual information. To train the encoders, we maximise the mutual information (MI) between the latent embeddings $z$ and the conditioning variables $\theta$. Algorithm 9 summarises the loss function.
Algorithm 9 InfoMax (Shannon-Jensen) Loss Computation